Grade 12Mechanics

two trains A &B are 100 Km apart traveling towards each other with starting speed for each 50 Km/hr trainA accelerate 20 & B retards 20 distance covered by A when they cross each other

divya
10 Years agoGrade 12

1 Answer

Abhay
8 Years ago
As s=ut+1/2at^2 here a is relative acc. i.e =0)So 100=100t+0 (here u relative equal to 50+50 as trains traveling in opposite direction)t=1So distance in 1 sec. by a equal ut+1/2at^2=50+1/2×20×1 (here a and u for a)=60km